A solution is prepared by dissolving 0.23 mol of hypchlorous and 0.27 mol of hypochlorite in water sufficient to yield 1.0 l of solution. The addition of 0.05 mol of HCI to this buffer solution caused the pH to drop slightly. The pH does not decrease drastically because the HCI reacts with the ____ present in the buffer solution. The Ka of hypchlorous a is is 1.36*10 to the -3
1
Expert's answer
2020-03-26T03:03:00-0400

pH of buffer solution can be calculated from acidity constant:

HClO <=> H+ + ClO-

[H+]*[ClO-]/[HClO]=Ka

[H+]=0.00136*0.23/0.27=0.00116 mole/l

pH=-lg[H+]=2.94

After addition of HCl it should shift the equilibrium in buffer solution:

HCl => H+ + Cl-

H+ + ClO- <=> HClO

the concentration of HClO would increase to 0.28 mole/l and concentration of ClO- should decrease to 0.22 mole/l, hence the pH changed to:

[H+]=0.00136*0.28/0.22=0.00173 mole/l

pH=-lg[H+]=2.76

This is the slight drop of pH from 2.94 to 2.76